OpenStudy (anonymous):

determine the focus for the equation y=x^2

OpenStudy (anonymous):

In y = ax^2, a = 1. But a = 1/(4p), where p is the distance from the vertex to the focus. So 1 = 1/4p 4p = 1 p = 1/4 The focus is 1/4 of a square up from the vertex inside the parabola at (0, 1/4). The directrix would then be 1/4 of a square below the vertex at y = -1/4.

OpenStudy (anonymous):

http://www.brightstorm.com/instantmath this website is great you type in any math problem and it nt only gives you the answers but the steps . :)